OpenStudy (anonymous):

What is the value of f(x) when x = -3 f(x)= -2x - 3 f(-3) = ?

OpenStudy (anonymous):

@michelle_DrNiah @nincompoop @baby1234567 @Vallarylynn16 @Callisto @Lena772 @kmacdaddy @JFraser @hlee13 @ganeshie8

OpenStudy (anonymous):

\[f(\spadesuit)=-2\times \spadesuit-3\] \[f(-3)=-2\times (-3)-3\]

OpenStudy (anonymous):

so is it -3 for the answer >

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this is the arithmetic problem you need to complete: \[-2\times (-3)-3\]

OpenStudy (anonymous):

first you need \(-2\times (-3)\) once you get that, subtract \(3\) from the result

OpenStudy (anonymous):

so its positve 3 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks can you help me with a couple more please

OpenStudy (anonymous):

Evaluate the function for x = -2a if a=3 f(x) =-4x + 7 @satellite73

OpenStudy (lena772):

still need help with that @aajkhdkhd? post a separate question & tag me! :)